Let A and B be subsets of U
n(U)=52
n(A)=25
n(B')=17
n(A U B)=41
Find n(B)?
It don't know where to start it

Answer by robertb(5830)

n(B) = n(U \ B') = n(U) - n(B') = 52 - 17 = 35.